OpenStudy (anonymous):

The correct answer to this is A, right? A polynomial function can be written as (x - 1)(x - 4)(x + 7). What are the x-intercepts of the graph of this function? a. (1, 0), (4, 0), (7, 0) b. (-1, 0), (-4, 0), (-7, 0) c. (1, 0), (4, 0), (-7, 0) d. (-1, 0), (-4, 0), (7, 0)

jimthompson5910 (jim_thompson5910):

(x - 1)(x - 4)(x + 7) = 0 x - 1 = 0 or x - 4 = 0 or x + 7 = 0 x = ??? or x = ??? or x = ???

jimthompson5910 (jim_thompson5910):

I'm using the zero product property for step 2

OpenStudy (anonymous):

So it would be C then?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

okay thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome